OpenStudy (anonymous):

Find the indefinite integral of: (X^2-4/x)dx

OpenStudy (anonymous):

seperate it: \[\frac{ x^2 -4 }{ x } = \frac{ x^2 }{ x } - \frac{ 4 }{ x } = x - \frac{ 4 }{ x }\] integrate seperately

OpenStudy (anonymous):

sepAration** lol

OpenStudy (anonymous):

is 4/x the answer

OpenStudy (anonymous):

i mean x-4/x

OpenStudy (isaiah.feynman):

Nope..

OpenStudy (anonymous):

no. integrate x - 4/x i simply separated it to make it integratable

OpenStudy (isaiah.feynman):

Integrable* lol